OpenStudy (anonymous):

WILL GIVE MEDALS open parentheses 7 minus square root of 54 close parentheses minus open parentheses 5 plus square root of 24 a) 2 minus square root of 78 B) 2 minus square root of 30 C) 2 minus 5 square root of 6 D) 2 minus square root of 6

OpenStudy (solomonzelman):

Lets write it out. \[(7-\sqrt{54})-(5+\sqrt{24})\]

OpenStudy (solomonzelman):

tell me if you lose me.\[(7-\sqrt{9 \times 6})-(5+\sqrt{4 \times 6})\]\[(7-3\sqrt{ 6})-(5+2\sqrt{ 6})\]

OpenStudy (anonymous):

ok whats next?

OpenStudy (solomonzelman):

the minus next to the second parenthesis, amkes both numbers in the second parenthesis negative. \[(7-3\sqrt{6})-(5+2\sqrt{6})~~~~~~~->~~~~~~7-3\sqrt{ 6}-5-2\sqrt{ 6}\]

OpenStudy (solomonzelman):

Add like terms, can you do that or should I ?

OpenStudy (anonymous):

i dont know how to combine the square roots, i know that 7-5 is 2

OpenStudy (anonymous):

is it five square root of 6?

OpenStudy (solomonzelman):

CLOSE \[2-5\sqrt{6}\]

OpenStudy (anonymous):

can u help me with another?

OpenStudy (anonymous):

5 fourth root of 7 end root minus 8 cubed root of 7 end root minus 2 fourth root of 7 end root minus cubed root of 7 a) 3 fourth root of 7 end root minus 9 cubed root of 7 b) Negative 6 square root of 14 c) Negative 6 square root of 7 d) 3 fourth root of 14 end root minus 9 cubed root of 14

OpenStudy (solomonzelman):

You subtract \[7-5~~~=2\] and \[-3\sqrt{6}-2\sqrt{6}=(-3-2)\sqrt{6}=-5\sqrt{6}\] for the question 1. so the final answer there would be \[2-5\sqrt{6}\]

OpenStudy (anonymous):

okay, i just forgot the 2 :)

OpenStudy (solomonzelman):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~Welcome~~to ~~~openstudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (solomonzelman):

I got disconnected, sorry.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 5\sqrt[4]{7}-8\sqrt[3]{7}-2\sqrt[4]{7} -\sqrt[3]{7}} \]

OpenStudy (solomonzelman):

I wrote it out, now lets see the like terms. \[\huge\color{green}{ 5\sqrt[4]{7}}\huge\color{red}{ -8\sqrt[3]{7}}\huge\color{green}{ -2\sqrt[4]{7}} \huge\color{red}{ -\sqrt[3]{7}} \]

OpenStudy (solomonzelman):

add greens and reds together.

OpenStudy (anonymous):

k so a) 3 fourth root of 7 end root minus 9 cubed root of 7

OpenStudy (solomonzelman):

YES!!!!!!!!

OpenStudy (anonymous):

Thanks :) im a fan :))

OpenStudy (solomonzelman):

\[^{yw,~~and~~TY}\]

OpenStudy (solomonzelman):

\(\huge\color{blue}{ (•‿•)} \)